Can you please help me with this problem? Using the function y=1/x-2
Give the y values for x= -2,-1,0,1,2,3.
Can you break it down and show me step by step?
Using these points, draw a curve and show it on a graph.

Answer by stanbon(75887)

y=1/x-2
y=1/(x-2)?
Give the y values for x= -2,-1,0,1,2,3.
If x=-2, y=1/(-4)=-1/4
If x=-1, y=1/(-3)=-1/3
If x=0, y=1/-2= -1/2
If x=2, y=1/0 which is undefined
If x=3, y=1/1=1
Cheers,
Stan H